# Technological Complexity Based on Japanese Patent Data

### v2

@@ Methods
The dataset used in this study consists of registered patents applied to the Japan Patent Office (JPO) over a 30-year period from fiscal 1981 to 2010. The data comprises detailed records of 3,189,536 patents filed by 64,622 corporations. Patent records follow the International Patent Classification (IPC) system and are further organized according to the Schmoch classification [28].
The Schmoch system aggregates IPC codes into 35 distinct technological fields, ensuring balanced class sizes and technological consistency. The Schmoch grouping has received widespread application in previous studies [8, 7, 10]. Regarding how to utilize a patent to measure contribution, traditional approaches have relied on raw patent counts to measure technological contributions in various fields [24, 7, 8, 9, 10, 33, 6]. However, as highlighted by Pintar et al. [8], simple counts may overestimate contributions when a patent involves multiple corporations or spans several technological fields. To address this issue, the primary IPC class of each patent is mapped onto the corresponding Schmoch field, and the value of a patent is fractionally allocated by dividing it equally among all associated corporations and among all technological fields to which it is attributed. This procedure prevents overestimation of contributions from a corporation or a field due to multiple co-ownership or cross-field classification. Although the raw dataset spans a wide range of patenting activity, most corporations show only minimal engagement such as holding a single patent in one field (SI Fig. S3). Low-activity cases may distort link formation in network analyses based on the HH algorithm [8]. For this reason, a filtering criterion based on the distribution of patent counts per corporation is therefore applied, retaining only the top 3% in terms of patent output. After filtering, the dataset consists of 2,485,027 patents from 1,938 corporations, representing approximately 93% of the original patent count.
@@
In order to quantify specialization in technological fields, computation of the revealed technological advantage (RTA)[29] index which builds on the revealed comparative advantage framework[30] is necessary. The RTA index for a given corporation in a specific technological field measures the ratio between the share of patents held by corporation $C$ in technology $T$ and the share of patents in technology $T$ among all corporations. Let $W_{C,T}$ denote the weighted patent count for corporation $C$ in technology $T$, where each patent is assigned a weight computed as one divided by the product of the number of associated corporations and the number of technological classes assigned to the patent. The numerator of the RTA index captures the fraction of patent activity devoted to technology $T$ for corporation $C$. Similarly, the denominator reflects the overall share of patent activity in technology $T$ among all corporations, computed as the sum of $W_{CT}$ over all corporations divided by the sum of $W_{CT}$ over all corporations and technologies. The RTA index is thus given by

### v3

@@ Methods
The dataset utilized in this study comprises registered patents submitted to the Japan Patent Office (JPO) covering fiscal years 1981 through 2010, encompassing detailed records of 3,189,536 patents filed by 64,622 corporations. Patent records adhere to the International Patent Classification (IPC) system and are further categorized according to the Schmoch classification [21].
The Schmoch system aggregates IPC codes into 35 distinct technological fields, ensuring balanced class sizes and technological consistency, and has been widely applied in previous research [29, 20, 30]. Nevertheless, the granularity of classification substantially influences analytical outcomes. Administrative classifications vary markedly in granularity and international comparability, potentially introducing analytical noise or biases if inadequately managed [8]. Employing consistent classification schemes, exemplified by the Schmoch system utilized in this study, mitigates such issues by providing balanced and coherent technological categories for robust analyses.
Traditional methods for measuring patent contributions typically employ raw patent counts [31, 20, 29, 32, 30, 17, 33]. However, such counting methods can lead to overestimation, particularly when patents involve multiple corporations or span multiple technological domains [29]. To address potential overestimations, the primary IPC class of each patent is mapped onto the corresponding Schmoch field, with patent value fractionally allocated by equally dividing it among associated corporations and technological fields. This fractional allocation prevents overestimation arising from multi-corporation ownership or cross-domain classifications.
Additionally, filtering criteria are essential to exclude extreme cases capable of distorting analytical outcomes [8, 29]. Accordingly, a filtering criterion based on patent distribution per corporation is applied in this study, retaining only the top 3% of corporations in patent output (Fig. A3). This filtering criterion excludes corporations exhibiting minimal patenting activity, thereby reducing significant analytical noise in network analyses. Following this filtering, the dataset comprises 2,485,027 patents filed by 1,938 corporations, representing approximately 93% of the original patent volume.
@@
Utilizing these filtered data, technological complexity is based on a bipartite network connecting corporations to technological fields in which they are specialized. Although there are some arguments regarding the optimal specialization index for establishing connections within these networks [29], we follows prior research by employing the Revealed Technological Advantage (RTA) index [34]. The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,the RTA index for a given corporation in a specific technological field measures the ratio between the share of patents held by corporation $C$ in technology $T$ and the share of patents in technology $T$ among all corporations. The numerator of the RTA index captures the fraction of patent activity devoted to technology $T$ for corporation $C$. Similarly, the denominator reflects the overall share of patent activity in technology $T$ among all corporations, computed as the sum of $W_{CT}$ over all corporations divided by the sum of $W_{CT}$ over all corporations and technologies. The RTA index is thus given by
@@
Data Availability
The patent data utilized in this study were obtained from the Japan Patent Office (JPO) via the J-platPat portal (https://www.j-platpat.inpit.go.jp/). Records published within the most recent twelve months are directly downloadable; data older than one year can be provided by the JPO upon reasonable request. Patent classification concordance tables mapping International Patent Classification (IPC) classes to Schmoch technology fields were acquired from the World Intellectual Property Organization (WIPO) IP Statistics portal (https://www.wipo.int/en/web/ip-statistics).
Code Availability
All scripts and code used to compute the Technological Complexity Index and reproduce the results of this paper are publicly available in the following GitHub repository: https://github.com/RintaroKARASHIMA/KCIinJapaneseFirms. The implementation of the Economic Complexity algorithms is based on the py-ecomplexity library (MIT License), available at https://github.com/cid-harvard/py-ecomplexity.